fr.uni Assignment ORION Chapter 04, Problem 014 (12.0 + (17.0p+ (6.50-and hen 770 s later has v-( 6 y + (170) + (16.0 k (in-pet second. (a)For that, 70s, A proton initially has in ti, has and then 7.70 s later has V4.60 +17.0-16.0k proto x axis? (c) Number Question Attempts: 0 of 8 used 4.24.12.1 422 PM A AE131/2019
Answer #1

a)

a = (v2 - v1) / t

a = [ (4 - 12) i + ( 17-17) j + ( - 16 - 6. 5)k] / 7.7

a = - 1.04 i + 0 j + 2.922 k

=======

b) magnitude

a^2 = 1.04^2 + 2.922^2

a= 3.1 m/s^2

======

c)

phi = 180 - arctan( 2.922/ 1.04)

phi = 109.6

======